OpenStudy (anonymous):

A chef is using a mixture of two brands of Italian dressing. The first brand contains 10% vinegar, and the second brand contains 13% vinegar. The chef wants to make 210 milliliters of a dressing that is 12% vinegar. How much of each brand should she use? (I've read like 4 different ways to do this and for some reason I can't wrap my head around it)

OpenStudy (kropot72):

Let the required quantity of the 10% vinegar dressing be a (ml), and let the required quantity of the 13% vinegar be b (ml). 0.1a + 0.13b = 210 * 0.12 ...................(1) a + b = 210 ................................(2) Now we can rearrange equation (2) to give b in terms of a as follows: b = 210 - a .................................(3) Next we can substitute the value of b given in equation (3) into equation (1) giving: 0.1a + 0.13(210 - a) = 210 * 0.12 which simplifies to: 0.1a + 27.3 - 0.13a = 25.2 .........(4) Now all that is needed is to solve equation (4) to find the value of a. Can you do that?

OpenStudy (anonymous):

I think my issue is that I'm not even understanding how everyone is getting from 1 to 2 on this. How do you get a+b=210 from 0.1a + 0.13b=210*0.12? Aren't the percentages need there?

OpenStudy (kropot72):

The percentages are put into decimal form in equation (1). 0.1a + 0.13b = 210 * 0.12 ...................(1) We have a (ml) of 10% vinegar dressing. Therefore the amount of vinegar in a (ml) of dressing is 0.1a. We have b (ml) of 13% vinegar dressing. Therefore the amount of vinegar in b (ml) of dressing is 0.13b We have 210 ml of 12% vinegar dressing. Therefore the amount of vinegar in 210 ml of the final dressing is 210 multiplied by 0.12. a + b = 210 ................................(2) We let 'a' be the required amount of 10% vinegar dressing and 'b' be the required amount of 13% vinegar dressing. Therefore when a and b are added they must make 210 ml.

OpenStudy (anonymous):

A=70, B=140?

OpenStudy (kropot72):

We have two unknown quantities in this question: 1. The required amount of 10% vinegar dressing. 2. The required amount of 13% vinegar dressing. Therefore we need to create a pair of simultaneous equations to be able to solve to find the values.

OpenStudy (anonymous):

am i even close..? lol

OpenStudy (kropot72):

Yes, your values for a and b are correct. 70 ml and 140 are the total amounts of the two original types of dressing. The amounts of vinegar in each are: 70 * 0.1 = 7 ml and 140 * 0.13 = 18.2 ml

OpenStudy (kropot72):

The amount of vinegar in the final dressing is: 210 * 0.12 = 25.2 ml which comes from: 7 + 18.2 = 25.2 ml